# WRF-GS: Wireless Radiation Field Reconstruction with 3D Gaussian Splatting

### v2

@@ Abstract
Wireless channel modeling plays a pivotal role in designing, analyzing, and optimizing wireless communication systems. Nevertheless, developing an effective channel modeling approach has been a longstanding challenge. This issue has been escalated due to the denser network deployment, larger antenna arrays, and wider bandwidth in 5G and beyond networks. To address this challenge, we put forth WRF-GS, a novel framework for channel modeling based on wireless radiation field (WRF) reconstruction using 3D Gaussian splatting. WRF-GS employs 3D Gaussian primitives and neural networks to capture the interactions between the environment and radio signals, enabling efficient WRF reconstruction and visualization of the propagation characteristics. The reconstructed WRF can then be used to synthesize the spatial spectrum for comprehensive wireless channel characterization. Notably, with a small number of measurements, WRF-GS can synthesize new spatial spectra within milliseconds for a given scene, thereby enabling latency-sensitive applications. Experimental results demonstrate that WRF-GS outperforms existing methods for spatial spectrum synthesis, such as ray tracing and other deep-learning approaches. Moreover, WRF-GS achieves superior performance in the channel state information prediction task, surpassing existing methods by a significant margin of more than 2.43 dB.
@@ I Introduction
Modern communications increasingly depend on wireless technologies that use electromagnetic (EM) waves for information exchange, driving advancements in mobile phones, automotive systems, and Internet-of-Thing devices [1]. At the core of these innovations lies wireless channel modeling, a long-standing problem in wireless communications. While the physics of EM wave propagation is described by Maxwell’s equations [2], solving these equations in reality is intricate due to the need for comprehensive knowledge of boundary conditions. This complexity has led to the development of various wireless channel modeling approaches, as shown in Fig. 1, which can be categorized into probabilistic modeling, deterministic modeling, and neural modeling.
@@
In contrast, the neural models, adopting the data-driven principle, learn the complex interactions between the environment and radio signals directly from location-based data, rather than relying on predefined statistical or physical models. One of the promising advancements in neural models is the recent adoption of the neural radiance field (NeRF), a breakthrough in computer vision for view synthesis [5]. Motivated by the fact that light is a kind of EM wave, Refs. [6] and [7] proposed two NeRF-based frameworks, named NeRF2 and NeWRF, respectively, for wireless channel modeling based on implicit wireless radiation field (WRF) reconstruction. The reconstructed WRF can provide detailed and accurate channel characteristics and significantly enhance communication performance. However, these methods suffer from high computational complexity and slow synthesis (a.k.a. rendering) speeds. In practice, the synthesis operation corresponds to the channel characteristic prediction process and will significantly affect the round-trip time in communication systems, which is critical for latency-sensitive applications. For instance, cloud gaming and digital twins typically demand low latency, i.e., under 20 milliseconds for cloud gaming [8] and often within a few milliseconds for digital twins [9]. This renders the NeRF-based methods impractical.
To conquer this challenge, we propose WRF-GS, a hybrid framework for fast and accurate channel modeling based on explicit WRF reconstruction using neural networks and 3D Gaussian splatting (3D-GS). 3D-GS utilizes an analytical architecture to represent and render 3D scenes, enabling high-quality and real-time view synthesis [10]. Compared with NeRF, 3D-GS offers faster rendering speeds and lower computational complexity. By adapting 3D-GS from the optical domain to the radio-frequency (RF) domain, WRF-GS can predict the received signal at any location within milliseconds for a give scene after training with a small number of signal measurements. Fig. 2 illustrates the synthesized spatial spectra of four algorithms based on the reconstructed WRF in a laboratory environment when the RX is fixed at one position and the TX is located at four different positions (i.e., P1-P4). Evidently, WRF-GS has the best prediction accuracy compared with the baselines, closely matching the ground truth.
@@
Figure 2: The synthesized spatial spectrum of different algorithms in a laboratory environment. The spatial spectrum represents the signal strength received from different directions composed of the azimuthal and elevation angles. (a) shows the 3D point clouds of the laboratory created by LiDAR, which are used for the ray-tracing algorithm. In addition, the TX can be located at any position, and the RX is equipped with a $4\times 4$ antenna array and fixed at a corner; (b) is a reconstructed WRF when the TX is located at P3 by running the WRF-GS framework using the 3D point clouds in (a) as initialization. The spatial spectrum is obtained by a series of nonlinear transformations, such as projecting the WRF onto the hemispherical plane, where the colors of the points represent different degrees of contribution to the spatial spectrum. (c) compares the synthesized spatial spectra of four algorithms when the TX is located at four different positions, i.e., P1-P4, as shown in (a). The ground truth is obtained using the antenna array.
@@
Nevertheless, directly applying the optical 3D-GS technique to the RF domain poses several challenges. First, the optical 3D-GS only considers the amplitude (i.e., light strength) of the optical signal, while the received RF signal consists of both amplitude and phase components, as EM waves are more prone to reflection, diffraction, and scattering than lights. Second, the measurement of visible light is via a million-pixel camera, but an RX typically utilizes either a single antenna or an antenna array. Third, the rendering function in the RF domain differs from the optical 3D-GS due to the unique physics of EM wave propagation. To address these challenges, we propose WRF-GS, a novel framework that refines the optical 3D-GS technique [10] from three key aspects. Scenario Representation Network: We introduce two multi-layer perceptions (MLPs) to effectively capture the complex interactions between the environment and radio signals. The learned features are then embedded into 3D Gaussian points, which serve as virtual TXs in wireless propagation. Projection Model: We devise a projection model to transform the optical camera model into an RF antenna model using spherical and Mercator projections. This model projects the virtual TXs onto the RX perception plane. Electromagnetic Splatting: Based on the projected 3D Gaussian points, we employ a differentiable tile rasterizer to synthesize wireless channel characteristics. This operation enables rapid synthesis of new channel characteristics within milliseconds.
@@
We conduct extensive experiments to evaluate the performance of the proposed WRF-GS. Unlike NeRF-based approaches that adopt an implicit representation, our WRF-GS explicitly represents the propagation characteristics, as shown in Fig. 2 (b). In addition, WRF-GS achieves higher reconstruction accuracy, higher sample efficiency, and faster rendering speed. Furthermore, we consider a downlink channel state information (CSI) prediction task in a multiple-input-multiple-output (MIMO) system as a case study to further demonstrate the effectiveness of WRF-GS in practical systems. Our results show that WRF-GS achieves better prediction performance of the downlink CSI than the baselines, outperforming NeRF2 by a significant margin of 2.43 dB.
@@
In summary, this paper makes the following contributions: i) We propose WRF-GS, a hybrid framework tailored for wireless channel modeling based on WRF reconstruction using neural networks and 3D-GS; ii) We adapt the optical 3D-GS to the RF domain by designing the modules of scenario representation network, projection model, and electromagnetic splatting. The integration of the EM wave physics and 3D-GS enables a visible, low-cost, and accurate channel modeling approach; iii) We conduct extensive experiments and a case study to demonstrate the effectiveness of WRF-GS. Numerical results show that WRF-GS outperforms existing methods.
@@ II Related Works
Environment-aware channel modeling has attracted much attention since future wireless networks will generate abundant location-specific channel data and are expected to possess more powerful data mining and AI capabilities [13, 14]. This motivated the adaptation of neural networks for more accurate channel modeling, i.e., learning the complex interactions between the environment and radio signals directly from location-based data [15]. To achieve this goal, a few recent works have been devoted to NeRF-based wireless channel modeling [16, 6, 7]. Orekondy et al. [16] proposed a neural surrogate to model wireless EM propagation effects in indoor environments based on NeRF using two synthetic datasets. Meanwhile, Zhao et al. [6] presented NeRF2 for spatial spectrum reconstruction in a real-world environment. By fixing a TX at an indoor location, Lu et al. [7] presented NeWRF, a NeRF-based wireless channel prediction framework with a sparse set of channel measurements. However, NeRF-based methods often suffer from high computational complexity and slower synthesis speed. In this paper, we present WRF-GS for environment-aware channel modeling with 3D-GS, reconciling the accuracy and computational complexity.
CSI prediction involves estimating future channel conditions based on current and historical data. Existing studies have been dedicated to predicting downlink channel states by observing uplink channels, as both are influenced by the same underlying physical environment and traverse identical paths [17, 18, 19]. Liu et al. [17] introduced FIRE, a variational auto-encoder (VAE) method, to transfer estimated CSI from the uplink to the downlink. In addition, Vasisht et al. [18] and Bakshi et al. [19] proposed R2F2 and OptML to predict the CSI using machine learning-based methods. In this paper, we apply WRF-GS to the CSI prediction task by transferring the estimated CSI from uplink to downlink.
@@ III Preliminaries
In this section, we present preliminaries on wireless channel modeling and 3D Gaussian Splatting.
@@ III-A Wireless Channel Modeling
A generic wireless communication system typically consists of a TX that generates and modulates an information signal, which is then transmitted through the wireless channel to an RX. The transmitted signal can be represented as a complex number $s=Ae^{j\varphi}$, where $A$ and $\varphi$ denote the amplitude and phase of the transmitted signal, respectively. In the wireless medium, the propagation of EM signals is characterized by key phenomena such as path loss and multipath propagation. For the basic free space path loss model, the received signal $y$ at the RX is given by
@@
where $\Delta A$ and $\Delta\varphi$ are the amplitude attenuation factor and phase rotation experienced by the signal, respectively. Due to various propagation effects, including reflection, scattering, refraction, and diffraction, the received signal is the sum of multiple copies of the transmitted signal. Consequently, it is more generally expressed as
@@
where $L$ is the total number of propagation paths, and $\Delta A_{l}$ and $\Delta\varphi_{l}$ represent the amplitude attenuation factor and phase rotation associated with the $l$-th path, respectively.
In multi-antenna systems, an antenna array is employed to characterize the energy distribution of the received signal $y$ from all directions, i.e., to estimate the AoA of the received signal. To illustrate this concept, we consider an antenna array equipped with $\sqrt{K}\times\sqrt{K}$ antennas, as depicted in Fig. 3. The spacing of two adjacent antennas is $D$, which is less than the wavelength $\lambda$. As shown in Fig. 3(b), the antenna array receives signals in a hemispherical plane to provide omnidirectional coverage and to efficiently capture signals from various directions. The direction of an RF source is characterized by the azimuthal angle $\alpha$ and elevation angle $\beta$, where $0^{\circ}\leq\alpha<360^{\circ}$ and $0^{\circ}\leq\beta<90^{\circ}$. The phase difference between the signals received by the $(m,n)$-th antenna pair is related to the distance of the two antennas from the source [20], i.e.,
@@
where $\Delta d_{m,n}$ is shown in Fig. 3(a) and can be computed by
@@
and $m,n=0,\ldots,\sqrt{K}-1$. In addition, $r_{m,n}=D\sqrt{m^{2}+n^{2}}$ and $\phi_{m,n}=\arctan 2(n,m)$ are the radius and angle of antenna $A_{m,n}$ at the polar coordinate system, respectively.
Figure 3: The AoA computation of the received signal with an antenna array. (a) There are $K$ antennas deployed in a grid, and the signal source is located at point B. (b) The spatial spectrum is generated from the antenna array. The illustration of spatial spectrum in Fig. 2(c) is obtained by projecting the spatial spectrum to the X-Y plane.
The AoA of the received signal is calculated by comparing the phases of signals received from multiple antennas. For AoA estimation, we use the antenna array to form a very narrow beam and steer it. When the beam is steered in the line-of-sight (LoS) direction, the array effectively blocks power from all other directions. This allows to reconstruct the spatial spectrum by steering the beam across a grid of azimuth and elevation angles and measuring the received signal power at each pointing direction. Specifically, we can form a matrix $\mathbf{P}$, where each element $\mathrm{P}(\alpha,\beta)$ represents the power of the received signal from the azimuthal angle $\alpha$ and elevation angle $\beta$, i.e.,
@@
where $\hat{\theta}_{m,n}$ denotes the measured phase of the RF signal received at the antenna $A_{m,n}$, and $\hat{\theta}_{0,0}$ is a constant term. Considering a one-degree angular resolution, the spatial spectrum can be reconstructed as a $360\times 90$ matrix:
@@
In summary, the reconstructed spatial spectrum can be interpreted as a spatial power distribution function, which directly corresponds to the WRF emitting from the transmitting source and propagating through the environment. In other words, the spatial spectrum serves as a representation of the underlying WRF, providing valuable insights into wireless channel modeling. Therefore, this paper aims to reconstruct the WRF and synthesize the spatial spectrum using the 3D-GS technique, which is introduced in the next subsection.
@@
3D-GS is a technique that utilizes a set of 3D Gaussian functions to model the geometry and appearance of a scene, offering a compact and efficient representation [10]. Unlike the computationally intensive and memory-heavy NeRF approach [5], 3D-GS provides a more lightweight and scalable solution for scene representation. Mathematically, 3D-GS provides a novel way to represent space using anisotropic Gaussian primitives, where each Gaussian is characterized by its own covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$ and center position vector $\boldsymbol{\mu}\in\mathbb{R}^{3}$. The 3D Gaussian distribution can be expressed as
@@
where $\boldsymbol{x}=(x_{0},x_{1},x_{2})$ represents the three-dimensional spatial position of the Gaussian point, and $\boldsymbol{\Sigma}$ can be expressed in terms of a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as
@@
The rendering process begins by transforming the 3D Gaussians $G(\boldsymbol{x})$ in space into 2D Gaussians $G^{\prime}(\boldsymbol{x^{\prime}})$ on the image plane [21]. These 2D Gaussians are then sorted based on their depth information. Each tile is processed independently and in parallel for the 2D Gaussians in its coverage. Utilizing the opacity ${o}_{i}$ and color attribute $c_{i}$ of each Gaussian, alpha blending is applied to the color values on each pixel:
$C=\displaystyle\sum_{i=1}^{N}{c}_{i}{\alpha}_{i}\displaystyle\prod_{j=1}^{i-1}(1-{\alpha}_{j})\ \mathrm{and}\ {\alpha}_{i}={o}_{i}{G}_{i}^{{}^{\prime}}(\varDelta\boldsymbol{p}_{i}),$ (7)
where $N$ is the number of 3D Gaussians near a given pixel, and $\varDelta\boldsymbol{p}_{i}=\boldsymbol{p}_{i}-\boldsymbol{p}_{s}$ is the difference vector between the projected center $\boldsymbol{p}_{i}$ of a Gaussian and the sampling pixel position $\boldsymbol{p}_{s}$. Once the color of each pixel is computed, the entire image is rendered. Subsequently, a comparison is made with the ground truth image to calculate the pixel-wise loss, which is then used to optimize the parameters of the model.

### v4

# Neural Representation for Wireless Radiation Field Reconstruction: A 3D Gaussian Splatting Approach

@@ Abstract
Wireless channel modeling plays a pivotal role in designing, analyzing, and optimizing wireless communication systems. Nevertheless, developing an effective channel modeling approach has been a long-standing challenge. This issue has been escalated due to denser network deployment, larger antenna arrays, and broader bandwidth in next-generation networks. To address this challenge, we put forth WRF-GS, a novel framework for channel modeling based on wireless radiation field (WRF) reconstruction using 3D Gaussian splatting (3D-GS). WRF-GS employs 3D Gaussian primitives and neural networks to capture the interactions between the environment and radio signals, enabling efficient WRF reconstruction and visualization of the propagation characteristics. The reconstructed WRF can then be used to synthesize the spatial spectrum for comprehensive wireless channel characterization. While WRF-GS demonstrates remarkable effectiveness, it faces limitations in capturing high-frequency signal variations caused by complex multipath effects. To overcome these limitations, we propose WRF-GS+, an enhanced framework that integrates electromagnetic wave physics into the neural network design. WRF-GS+ leverages deformable 3D Gaussians to model both static and dynamic components of the WRF, significantly improving its ability to characterize signal variations. In addition, WRF-GS+ enhances the splatting process by simplifying the 3D-GS modeling process and improving computational efficiency. Experimental results demonstrate that both WRF-GS and WRF-GS+ outperform baselines for spatial spectrum synthesis, including ray tracing and other deep-learning approaches. Notably, WRF-GS+ achieves state-of-the-art performance in the received signal strength indication (RSSI) and channel state information (CSI) prediction tasks, surpassing existing methods by more than 0.7 dB and 3.36 dB, respectively. The code is available at https://github.com/wenchaozheng/WRF-GSplus.
@@ I Introduction
Modern communications increasingly depend on wireless technologies that use electromagnetic waves (EM) for information exchange, driving advances in mobile phones, automotive systems, and the Internet of Things [2]. At the core of these advances lies wireless channel modeling, a long-standing problem for modeling the interactions between the environment and radio signals in wireless communications. While Maxwell’s equations model such interactions [3], solving these equations in reality is intricate due to the need for comprehensive knowledge of boundary conditions. This complexity has led to the development of various wireless channel modeling approaches, as shown in Fig. 1, which are categorized into probabilistic modeling, deterministic modeling, and neural modeling.
@@
In contrast, the neural models, adopting the data-driven principle, learn the complex interactions between the environment and radio signals directly from location-based data. One of the promising advancements in neural models is the recent adoption of the neural radiance field (NeRF), a breakthrough in computer vision for view synthesis [6]. Motivated by the fact that light is a kind of EM wave, Refs. [7] and [8] proposed two NeRF-based frameworks, named NeRF2 and NeWRF, respectively, for wireless channel modeling based on implicit wireless radiation field (WRF) reconstruction. The reconstructed WRF can provide detailed and accurate channel characteristics and significantly enhance communication performance. However, these methods suffer from high computational complexity and slow synthesis (a.k.a. rendering) speeds. In practice, the synthesis operation corresponds to the channel characteristic prediction process and will significantly affect the round-trip time in communication systems, which is critical for latency-sensitive applications. For instance, cloud gaming and digital twins typically demand low latency, i.e., under 20 milliseconds for cloud gaming [9] and often within a few milliseconds for digital twins [10]. This renders the NeRF-based methods impractical.
@@
Figure 2: The synthesized spatial spectra of five algorithms in a laboratory environment. The spatial spectrum represents the signal strength received from different directions composed of the azimuthal and elevation angles. (a) shows the 3D point clouds of the laboratory created by LiDAR, which are used for the ray tracing algorithm. In this setting, the TX can be located at any position, while the RX is equipped with a $4\times 4$ antenna array and fixed at a corner. (b) compares the synthesized spatial spectra of five algorithms when the TX is located at four different positions, i.e., P1-P4, as shown in (a). The ground truth is obtained using the antenna array.
Recently, 3D Gaussian splatting (3D-GS) has emerged as a promising technique for efficient and compact scene reconstruction using a set of 3D Gaussian functions [11]. Compared with NeRF, 3D-GS offers significantly faster rendering speeds and lower computational complexity. These advantages motivate us to adapt the 3D-GS technique for reconstructing the WRF. Nevertheless, directly applying the optical 3D-GS to the RF domain poses several significant challenges. First, the optical 3D-GS only considers the amplitude of optical signals (i.e., light intensity). In contrast, the received RF signals consist of both amplitude and phase components, as EM waves are more prone to reflection, diffraction, and scattering than visible light. Second, visible light is typically measured using million-pixel cameras, while an RX employs either a single antenna or an antenna array. Third, the rendering function in the RF domain differs from that in optical 3D-GS due to the unique physics of EM wave propagation.
@@
To address these challenges, we propose WRF-GS, a novel framework for fast and accurate channel modeling based on explicit WRF reconstruction using neural networks and 3D-GS. Specifically, we refine the optical 3D-GS technique from three key aspects to adapt it to the RF domain. First, we design a scenario representation network that comprises two multilayer perceptions (MLPs) to effectively capture the complex interactions between the environment and radio signals. The learned features are then embedded into 3D Gaussian points, which serve as virtual TXs to model wireless signal propagation. This approach enables the network to represent the spatial distribution of signals. Second, we devise a projection model to adapt the optical camera model to an RF antenna model using spherical and Mercator projections. This model projects the virtual TXs onto the RX perception plane, enabling accurate spatial mapping of signal propagation. Third, we introduce an electromagnetic splatting process using a differentiable tile rasterizer to synthesize wireless channel characteristics based on the projected 3D Gaussian points. This process, combined with the projection model, enables rapid synthesis of new channel characteristics within milliseconds.
However, WRF-GS faces limitations in accurately modeling high-frequency signal variations and also suffers from parameter redundancy due to its coarse neural network design. To overcome these limitations, we further put forth WRF-GS+, an enhanced version of WRF-GS that integrates electromagnetic wave physics into the neural network design and eliminates the need for additional parameters. Specifically, we adopt the deformable 3D Gaussians in the scenario representation network to model both static and dynamic components of the WRF, significantly improving the ability to capture the complex interactions between the environment and radio signals. In addition, we exploit the inherent properties of 3D Gaussian primitives directly and leverage the $\alpha$-blending technique to synthesize wireless channel characteristics based on the projected 3D Gaussian points. These enhancements enable WRF-GS+ to characterize the distribution of the WRF in a given environment more effectively and with greater computational efficiency. Fig. 2 illustrates the synthesized spatial spectra of five algorithms based on the reconstructed WRF in a laboratory environment. Evidently, WRF-GS+ achieves the best prediction accuracy compared with the WRF-GS method and other baselines, closely matching the ground truth.
@@
We conduct extensive experiments to evaluate the WRF-GS and WRF-GS+ methods. The numerical results show that the proposed methods are capable of explicitly representing the propagation characteristics and achieving higher reconstruction accuracy, greater sample efficiency, and faster rendering speed in wireless channel modeling. Moreover, we present two case studies, i.e., the received signal strength indicator (RSSI) prediction and the downlink channel state information (CSI) prediction, to demonstrate the effectiveness of the proposed methods in practical systems. The RSSI prediction task focuses on estimating the signal strength at a specific location, while the CSI prediction task involves estimating the downlink CSI based on the uplink CSI measurements in a multiple-input-multiple-output (MIMO) system. The numerical results show that WRF-GS+ achieves superior performance in both tasks, outperforming existing methods by more than $0.7$ dB in RSSI prediction and $3.36$ dB in CSI prediction.
Our main contributions are summarized as follows:
  • We propose WRF-GS, a hybrid framework tailored for wireless channel modeling based on WRF reconstruction using neural networks and 3D-GS. This framework enables efficient and accurate modeling of wireless propagation characteristics.
@@
  • We adapt the optical 3D-GS to the RF domain by designing the modules of scenario representation network, projection model, and electromagnetic splatting. The integration of the EM wave physics and 3D-GS enables a visible, low-cost, and accurate channel modeling.
  • We propose WRF-GS+, an enhanced version of WRF-GS, which employs deformable 3D Gaussians to separately model the static and dynamic components of the signals from virtual TXs. This approach improves efficiency and accuracy in characterizing the distribution of WRF.
  • We conduct extensive experiments and two case studies to demonstrate the effectiveness of the WRF-GS and WRF-GS+ methods. Numerical results show that the proposed methods consistently outperform the existing methods.
The remainder of this paper is organized as follows. Section II reviews the related works. Section III provides the preliminaries of the wireless channel model and 3D-GS. Sections IV and V detail the proposed WRF-GS and WRF-GS+ methods, respectively. Section VI presents the implementation and evaluation of the proposed methods in a laboratory environment using a real-world dataset. Two case studies are presented in Section VII to further demonstrate the practical applicability of the methods. Finally, Section VIII concludes the paper.
@@ II Related Works
Environment-aware channel modeling has attracted much attention since future wireless networks will generate abundant location-specific channel data and are expected to possess more powerful data mining and AI capabilities [14, 15]. This motivated the adaptation of neural networks for more accurate channel modeling, i.e., learning the complex interactions between the environment and radio signals directly from location-based data [16, 17]. To achieve this goal, a few recent works have been devoted to NeRF-based wireless channel modeling [18, 7, 8]. Orekondy et al. [18] proposed a neural surrogate to model wireless EM propagation effects in indoor environments based on NeRF using two synthetic datasets. Meanwhile, Zhao et al. [7] presented NeRF2 for spatial spectrum reconstruction in a real-world environment. By fixing a TX at an indoor location, Lu et al. [8] presented NeWRF, a NeRF-based wireless channel prediction framework with a sparse set of channel measurements. However, NeRF-based methods often suffer from high computational complexity and slower synthesis speed. In this paper, we present WRF-GS for environment-aware channel modeling with 3D-GS, reconciling the accuracy and computational complexity.
RSSI prediction is a typical task in wireless networks to estimate the signal strength at a specific location based on factors such as distance, environment, and prior measurements. This task is crucial for network management, indoor positioning, and signal coverage planning. Shin et al. [19] proposed an autonomous smartphone-based war-walking system (MRI) that automatically constructs a radio fingerprint database using limited user-collected data. MRI uses a log-distance path loss model to predict RSSI values throughout the entire indoor space based on partial measurements. Similarly, Zhao et al. [7] devised a turbo learning-enabled RSSI prediction method based on the NeRF2 framework using the Bluetooth Low Energy (BLE) dataset. In this paper, we apply the WRF-GS and WRF-GS+ methods to the RSSI prediction task by retraining them on the public BLE dataset [7].
CSI prediction involves estimating future channel conditions based on current and historical data. Existing studies have focused on predicting downlink channel states by observing uplink channels, as both are influenced by the same underlying physical environment and traverse identical paths [20, 21, 22]. Liu et al. [20] introduced FIRE, a variational auto-encoder (VAE) method, to transfer estimated CSI from the uplink to the downlink. In addition, Vasisht et al. [21] and Bakshi et al. [22] proposed R2F2 and OptML to predict the CSI using machine learning-based methods. In this paper, we apply the WRF-GS and WRF-GS+ methods to the CSI prediction task by transferring the estimated CSI from uplink to downlink.
@@ III Preliminaries
In this section, we present preliminaries on wireless channel modeling and 3D-GS technique.
@@ III-A Wireless Channel Modeling
A generic wireless communication system typically consists of a TX that generates and modulates an information signal, which is then propagated through the wireless channel to an RX. The transmitted signal can be mathematically represented as a complex number $s=Ae^{j\varphi}$, where $A$ and $\varphi$ denote the amplitude and phase of the signal, respectively. In the wireless medium, the propagation of EM signals is influenced by fundamental physical phenomena, including path loss and multipath propagation. In the case of free space path loss, the received signal $y$ at the RX is expressed as:
@@
where $\Delta A$ and $\Delta\varphi$ are the amplitude attenuation factor and phase rotation incurred during propagation, respectively. However, in realistic environments, the signal encounters a multitude of propagation effects, including reflection, scattering, refraction, and diffraction. These phenomena cause the transmitted signal to split into multiple replicas, each traversing different paths before arriving at the RX. As a result, the received signal is more accurately modeled as the superposition of these multipath components:
@@
where $L$ denotes the total number of propagation paths, and $\Delta A_{l}$ and $\Delta\varphi_{l}$ correspond to the amplitude attenuation factor and phase rotation specific to the $l$-th path, respectively. This multipath formulation captures the complex interplay of signal interactions with the environment, providing a comprehensive framework to characterize the behaviors of wireless channels.
In multi-antenna systems, an antenna array is employed to characterize the spatial energy distribution of the received signal $y$ from all directions, i.e., to estimate the AoA. To illustrate this concept, we consider an antenna array equipped with $\sqrt{K}\times\sqrt{K}$ antennas, as depicted in Fig. 3. The spacing of two adjacent antennas is $D$, where $D<\lambda$ and $\lambda$ denotes the wavelength of the transmitted signal. As shown in Fig. 3(b), the antenna array operates within a hemispherical plane, providing omnidirectional coverage to efficiently capture signals from various directions. The direction of an RF source is characterized by two angles: the azimuthal angle $\alpha$ $(0^{\circ}\leq\alpha<360^{\circ})$ and the elevation angle $\beta$ $(0^{\circ}\leq\beta<90^{\circ})$. The phase difference between the signals received by the $(m,n)$-th antenna pair is determined by the relative distance of the two antennas from the source [23], expressed as:
@@
where $\Delta d_{m,n}$, illustrated in Fig. 3(a), is computed as
@@
Here, $m,n=0,\ldots,\sqrt{K}-1$, $r_{m,n}=D\sqrt{m^{2}+n^{2}}$ represents the radial distance of antenna $A_{m,n}$ in a polar coordinate system, and $\phi_{m,n}=\arctan 2(n,m)$ is its angular position.
Figure 3: The AoA computation of the received signal with an antenna array. (a) There are $K$ antennas deployed in a grid, and the signal source is located at point B. (b) The spatial spectrum is generated from the antenna array. The illustration of the spatial spectrum in Fig. 2(c) is obtained by projecting the spatial spectrum to the X-Y plane.
The AoA of the received signal is calculated by comparing the phase differences of the signals received across multiple antennas. To achieve this, we use the antenna array to form a narrow beam and steer it across different directions. When the beam aligns with the line-of-sight (LoS) direction, the array effectively suppresses signals from other directions. This allows us to reconstruct the spatial spectrum by steering the beam across a grid of azimuth and elevation angles and measuring the received signal power at each pointing direction. Specifically, we can form a matrix $\mathbf{P}$, where each element $\mathrm{P}(\alpha,\beta)$ represents the received signal power from the azimuthal angle $\alpha$ and elevation angle $\beta$, i.e.,
@@
where $\hat{\theta}_{m,n}$ denotes the measured phase of the RF signal received at antenna $A_{m,n}$, and $\hat{\theta}_{0,0}$ is a constant reference phase. Considering a one-degree angular resolution, the spatial spectrum is reconstructed as a $360\times 90$ matrix:
@@
The reconstructed spatial spectrum represents a spatial power distribution function, which directly corresponds to the WRF emitted by the transmitting source and propagated through the environment. In other words, the spatial spectrum serves as a representation of the underlying WRF, providing valuable insights into wireless channel modeling. Therefore, this paper aims to reconstruct the WRF and synthesize the spatial spectrum using the 3D-GS technique, which is introduced in the following subsection.
@@
3D-GS is a technique that utilizes a set of 3D Gaussian functions to model the geometry and appearance of a scene, offering a compact and efficient representation [11, 24]. Mathematically, 3D-GS represents space using anisotropic Gaussian primitives, where each Gaussian is characterized by its covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$ and center position vector $\boldsymbol{\mu}\in\mathbb{R}^{3}$. The 3D Gaussian distribution is expressed as
@@
where $\boldsymbol{x}=(x_{0},x_{1},x_{2})$ represents the spatial coordinates of a Gaussian point in 3D space, and $\boldsymbol{\Sigma}$ can be expressed in terms of a scaling matrix $\mathbf{S}$ and a rotation matrix $\mathbf{R}$ as
@@
The rendering process begins by projecting the 3D Gaussians $G(\boldsymbol{x})$ onto the image plane, resulting in 2D Gaussians $G^{\prime}(\boldsymbol{x^{\prime}})$ [25]. These 2D Gaussians are then sorted based on their depth information to ensure proper occlusion handling. Each tile on the image plane is processed independently and in parallel for the 2D Gaussians within its coverage. Using the opacity ${o}_{i}$ and color attribute $c_{i}$ of each Gaussian, $\alpha$-blending is applied to compute the color values for each pixel:
$C=\displaystyle\sum_{i=1}^{N}{c}_{i}{\alpha}_{i}\displaystyle\prod_{j=1}^{i-1}(1-{\alpha}_{j}),$ (7)
where
${\alpha}_{i}={o}_{i}{G}_{i}^{{}^{\prime}}(\varDelta\boldsymbol{p}_{i}).$
Here, $N$ denotes the number of 3D Gaussians affecting a given pixel, and $\varDelta\boldsymbol{p}_{i}=\boldsymbol{p}_{i}-\boldsymbol{p}_{s}$ represents the difference vector between the projected center $\boldsymbol{p}_{i}$ of a Gaussian and the sampling pixel position $\boldsymbol{p}_{s}$. Once the color of each pixel is computed, the entire image is rendered. The rendered image is then compared with the ground truth to calculate the pixel-wise loss, which is used to optimize the model parameters.
In summary, 3D-GS is an efficient and scalable technique for representing and rendering 3D scenes in computer graphics by modeling scenes as a collection of Gaussian primitives. This motivates us to adopt the 3D-GS technique for wireless channel modeling based on explicit WRF reconstruction. This adaptation is achieved by treating Gaussian primitives as virtual TXs that capture wireless signal strengths and attenuation properties. By mapping the learned Gaussian representations onto the receiver’s hemispherical antenna perception plane, we can efficiently reconstruct the WRF and rapidly synthesize accurate spatial power spectra in Eqn. (4) and Fig. 2(b).